OpenStudy (anonymous):

find the new equation if the origin is (1,2) 5x+4y+3=0

OpenStudy (anonymous):

This is a translation to the left by 1 and down by 2 (since the origin moved the opposite directions...right 1 and up 2). Plug (into your original equation) x + 1 for x and y + 2 in for y and simplify.

OpenStudy (anonymous):

(5x-1)+(4x-2)+3=0 -5-8+3=0

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

Thats wrong im an idiot lol

OpenStudy (anonymous):

@accqqk13 I've got the same answer like you, is that correct @mtbender74?

OpenStudy (anonymous):

No...@aCQQK13's answer is not correct. All they did was plug in the new origin coordinates. Try graphing it and then imagine if you slide the origin to the new location (without moving the graph... Moving the origin is the same as translating the graph. THe direction the origin moves is opposite the direction the graph moves. At that point, you just need to apply your translation rules...which is what I did in my post.

OpenStudy (anonymous):

no my answer = -10

OpenStudy (anonymous):

-5-8+3=-10 -5-8+3+10=0

OpenStudy (anonymous):

do i have to use completing the square method in this equation ?

OpenStudy (anonymous):

no...there should be no completing the square...that is for quadratics. This is finding the new equation for a line that has been translated.

OpenStudy (anonymous):

im confuse, can you give an example problem regarding this? thank you really appreciate it .: )

OpenStudy (anonymous):

Ok...imagine you have the line y=x. This line has a slope of 1 and goes through the point (0,0). If you shift the origin to (1,2), this line will have the same slope as before (m=1) but will now go through the point (-1,-2). With me so far?

OpenStudy (anonymous):

oh this was a translation - i couldn't tell

OpenStudy (anonymous):

yeah,, i got the i dea.

OpenStudy (anonymous):

*idea

OpenStudy (anonymous):

So to find the new equation, we substitute x and y from the original with x + 1 and y + 2.

OpenStudy (anonymous):

(5x-1)+(4x-2)+3=0 ?

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

It should be 5(x+1) + 4(x+2) + 3 = 0

OpenStudy (anonymous):

You replace the x with x +1 and the y with y + 2...understand?

OpenStudy (anonymous):

Now, you just distribute and simplify...

OpenStudy (anonymous):

FYI with mtbender74's 5x+5+4x+8+3=0 9x+16=0

OpenStudy (anonymous):

5(x+1) + 4(y+2) +3 = 0 5x +5 + 4y + 8 + 3 = 0 5x + 4y + 16 = 0 (this is the answer) I think sir mtbender74 just forgot it's y from his equation "5(x+1) + 4(x+2) + 3 = 0 for 4(x+2)" It should be 4(y+2)